A flywheel has a moment of inertia of 1.6 × 10−3 kg.m2. When a constant torque is applied, it reaches an angular velocity of 1200 rev/min in 15 s. Assuming it started from rest, find:
i) the angular acceleration
ii) the unbalanced torque applied
iii) the angle turned through 15 s
iv) the work done on the flywheel by the torque
1
Expert's answer
2020-05-26T12:49:05-0400

Given, Moment of inertia "I=1.6 \u00d7 10^{\u22123} kg.m^2" ,

Initial angular speed "\omega_0=0" ,final angular speed "\omega=1200\times \frac{2\pi}{60}=40\pi \:rad/s" ,time "t=15s" ,

i). Angular acceleration

"\alpha=\frac{\omega-\omega_0}{t}\implies \alpha=\frac{40\pi}{15}=\frac{8\pi}{3} \: rad/s^2"

ii).

Unbalanced torque

"\tau=I\alpha\\
\implies \tau=1.6\times 10^{-3}\times \frac{8\pi}{3}=13.40\times 10^{-3}\:Nm"

iii). Let the required angle is "\theta" , thus

"\omega^2=\omega_0^2+2\alpha \theta\\
\implies \theta=\frac{\omega^2-\omega_0^2}{2\alpha}\\
\implies \theta=\frac{(40\pi)^2}{2\frac{8\pi}{3}}=942.47 \: rad"

Alternatively, we can apply

"\theta=\omega_0t+\frac{1}{2}\alpha t^2"

and eventually we will get the same result.



iv).

By applying work energy theorem for rotational motion we get,


"W_{rot}=\frac{1}{2}I\omega^2-\frac{1}{2}I\omega_0^2"

Thus we get,

"W_{rot}=\frac{1}{2}\times 1.6\times 10^{-3} (40\pi)^2=12.63J"

Hence we done.